# A Bayesian approach for fitting semi-Markov mixture models of cancer latency to individual-level data

### v1

@@ 4 Markov chain Monte Carlo scheme
Next, we briefly review Gibbs and Metropolis-Hastings updates (Brooks et al., 2011). The Gibbs step for $\beta$ proceeds by drawing a new value for this parameter from its full conditional distribution (19) given the most recent values of $\mathbf{Z}$. For the Metropolis-Hasting steps, consider the random vector $(\zeta,\eta)$ where $\zeta$ and $\eta$ are random vectors and from which we wish to update $\zeta$. Given current values $(\tilde{\zeta},\tilde{\eta})$, a candidate value $\zeta^{\star}$ is drawn according to the proposal density $q(\zeta\mid\tilde{\zeta},\tilde{\eta})$, and is accepted with probability
$\alpha\left(\tilde{\zeta},\tilde{\eta}\rightarrow\zeta^{\star},\tilde{\eta}\right)=\min\left\{1,\dfrac{\pi\left(\zeta^{\star},\tilde{\eta}\right)q\left(\tilde{\zeta}\mid\zeta^{\star},\tilde{\eta}\right)}{\pi\left(\zeta^{(m)},\tilde{\eta}\right)q\left(\zeta^{\star}\mid\tilde{\zeta},\tilde{\eta}\right)}\right\}=\min\left\{1,\dfrac{\pi\left(\zeta^{\star}\mid\tilde{\eta}\right)q\left(\tilde{\zeta}\mid\zeta^{\star},\tilde{\eta}\right)}{\pi\left(\zeta^{(m)}\mid\tilde{\eta}\right)q\left(\zeta^{\star}\mid\tilde{\zeta},\tilde{\eta}\right)}\right\},$ (20)
where $\pi(\eta,\zeta)$ is the target distribution; otherwise, the current value is kept. The proposal distributions for $Z^{HP}_{i}$ and $\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n}\right)$ are described in Sections 4.1 are 4.2 respectively. Naturally, the proposal distributions for $\theta$ depend on the choice of $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$.
4.1 Metropolis-Hastings update for $Z^{HP}_{i}$
The efficiency of the Metropolis-Hastings steps for $Z^{HP}_{1},\dots,Z^{HP}_{n}$ is central to the overall efficiency of our data-augmented Markov chain Monte Carlo sampler. We construct a proposal distribution for $Z^{HP}_{i}$ that is independent of the current value of this latent variable; a higher acceptance rate, therefore, results in better mixing. To this end, we design a proposal distribution that is close to the full conditional $\pi\left(Z^{HP}_{i}\mid\cdot\right)$. The support of $\pi\left(Z^{HP}_{i}\mid\cdot\right)$ is $(t_{0},c_{i}]\cup\varnothing$ for individuals in the censored group and $(t_{0},c_{i}]$ for those in the screen-detected and interval-detected groups. We therefore consider a proposal distribution with a positive density on the interval $(t_{0},c_{i}]$ and a point mass at $\varnothing$ for the censored group. The exact form of the proposal distribution depends on the group of the individual, described below.
4.1.1 Censored group
For an individual $i$ in the censored group, the observed data consist of $n_{i}$ negative screens and $t^{PC}_{i}=c_{i}$. The support of the full conditional distribution of $Z^{HP}_{i}$ is $(t_{0},c_{i}]\cup\varnothing$, and from Equations (9), (11) and (14), (15), its density is
$\pi\left(Z^{HP}_{i}\mid\cdot\right)\propto L_{i}(\theta,\psi;Z^{HP}_{i},Z^{I}_{i})f_{t}\left(t^{PC}_{i}\mid Z^{HP}_{i},Z^{I}_{i},\theta\right)f_{O}\left(O^{i}_{1},\dots,O^{i}_{n_{i}}\mid Z^{HP}_{i},\beta\right)$
$\quad\propto\left[\delta_{\varnothing}\left(Z^{HP}_{i}\right)S_{\textsf{H}}(c_{i}-t_{0})+\delta_{(t_{0},c_{i})}\left(Z^{HP}_{i}\right)f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)^{1-Z^{I}_{i}}\right]\beta^{m_{i}^{+}}(1-\beta)^{m_{i}^{-}\left(Z^{HP}_{i}\right)}$
$\quad=\delta_{\varnothing}\left(Z^{HP}_{i}\right)S_{\textsf{H}}(c_{i}-t_{0})+\delta_{(t_{0},c_{i})}\left(Z^{HP}_{i}\right)f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)^{1-Z^{I}_{i}}(1-\beta)^{m_{i}^{-}\left(Z^{HP}_{i}\right)}.$ (21)
Here $m_{i}^{+}=0$ because person $i$ belongs to the censored group, the notation $m_{i}^{-}\left(Z^{HP}_{i}\right)$ emphasizes that the number of false negative screens is a function of $Z^{HP}_{i}$, and the equality holds because $m_{i}^{-}\left(\varnothing\right)=0$. The full conditional (4.1.1) has a point mass at $\varnothing$ and a positive density on $(t_{0},c_{i}]$. The integer-valued quantity $m^{-}\left(Z^{HP}_{i}\right)$ induces discontinuities, as shown in Figure 3(a).
We construct a proposal distribution for the Metropolis-Hastings step that mimics the full conditional  (4.1.1). First, the proposal has a point mass component at $\varnothing$ with weight proportional to $S_{\textsf{H}}(c_{i}-t_{0})$. The density component is constructed as a mixture distribution with a discontinuous density proportional to $f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)(1-\beta)^{m_{i}^{-}\left(Z^{HP}_{i}\right)}$. The expression is designed to be equal to the full conditional distribution whenever the pre-clinical cancer is indolent ($Z^{I}_{i}=1$), and otherwise only differs by the factor $S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)$ for the case of progressive pre-clinical cancer. The mixture components are truncated Weibull distributions restricted to non-overlapping supports defined by the inter-screen intervals. Over the $k$th inter-screen interval $(t^{i}_{k-1},t^{i}_{k}]$, the truncated Weibull density of the corresponding mixture component is proportional to $f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)$, and its mixture weight is proportional to

### v2

@@ 4 Markov chain Monte Carlo scheme
Algorithm 1 Steps of the DA-Markov chain Monte Carlo sampler.
4.1 Theoretical guarantees
Theorem 4.1 shows that the Markov chain underlying the DA-MCMC sampler is ergodic; its proof appears in web Appendix C. The ergodic theorem, therefore, holds for any $\pi$-integrable function $h$, and the estimator $\bar{h}_{M}=\frac{1}{M}\sum_{m=1}^{M}h\left(\theta^{(m)},\psi^{(m)},\beta^{(m)},\mathbf{Z}^{(m)}\right)$ is asymptotically consistent for $E_{\pi}h$ under any initial distribution 28.
Theorem 4.1 (Ergodicity).
The Markov chain $\left\{\left(\theta^{(m)},\psi^{(m)}\beta^{(m)},\mathbf{Z}^{(m)}\right)\right\}_{m=1}^{M}$ underlying the DA-MCMC algorithm is ergodic.
4.2 Model comparison
Mis-specifying a model of cancer natural history can lead to biased inference 17. To identify specifications for $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$ that fit the data well, we compare the predictive fit of models with different choices for these distributions. We accomplish this by extending ALOOCV 20, 21 to the class of latent variable models. While exact CV requires fitting the model multiple times, approximate CV can be easily computed from the posterior draws of a MCMC sampler. Once we have obtained predictive fit estimates for each model under consideration, we remove the models whose predictive fit is significantly lower than that of the best model, as determined by pairwise t-tests. We then proceed by choosing among the remaining models the one whose specification for $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$ we deem agrees the most with existing medical knowledge. The procedure is straightforward to implement; web Appendix F contains the complete details of our extension of ALOOCV to latent models.
4.3 Model implementation
The methodology described in Section 3.3 applies to any continuous distribution $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$ on the positive line. For the simulations in Section 5 and the case study in Section 6, we choose $\textsf{F}_{\textsf{H}}=\textsf{Wei}(\lambda_{\textsf{H}},\alpha_{\textsf{H}})$ and $\textsf{F}_{\textsf{prog}}=\textsf{Wei}(\lambda_{\textsf{prog}},\alpha_{\textsf{prog}})$, with $\textsf{Wei}(\lambda,\alpha)$ denoting the Weibull distribution with rate $\lambda>0$ and shape $\alpha>0$, and whose density is proportional to $x^{\alpha-1}\exp\{-\lambda x^{\alpha}\}$. The Weibull choice has several advantages, including flexibility and interpretability: if $\alpha>1$ ($\alpha<1)$, the hazard function increases (decreases) with age as $\propto t^{\alpha-1}$; and if $\alpha=1$ it corresponds to the exponential distribution. In addition, the Weibull distribution’s survival function has a closed-form expression, and its cumulative distribution function can be inverted, enabling efficient sampling from $\textsf{Trunc-F}_{\textsf{H}}$ in the proposal; distribution described in web Appendix B.‘
We implement the MCMC sampler in the programming language C++ for efficiency. The sampler is fast, scaling to cohort studies with tens of thousands of individuals. For convenience, the C++ implementation is made available in R with the help the package RCPP 29 via the R package baclava available on CRAN. In the MCMC runs in Sections 5-6, we keep $10^{3}$ thinned draws post warm-up phase due to memory constraints.
5 Performance on simulated data
To examine the mixing properties of our sampler, we run multiple MCMC samplers initialized at over-dispersed values on a simulated dataset. We then use the potential scale reduction factor 30, 31, the effective sample size, and traceplots to assess the algorithm’s mixing properties.
For the purposes of a simulated dataset, we set the starting age for non-zero onset hazard to $t_{0}=30$ years, and assume a linear increase in both the onset and progression hazards ($\alpha_{H}=\alpha_{prog}=2$). We further set the Weibull rates to $\lambda_{\textsf{H}}=6.5\times 10^{-5}$ and $\lambda_{\textsf{prog}}=3.14\times 10^{-2}$, resulting in a cumulative risk of pre-clinical onset by age $80$ of $15\%$, and a mean sojourn time of $\mu_{\textsf{prog}}=5$ years. Finally, we set the indolent probability to $\psi=0.1$ and the screen sensitivity to $\beta=0.85$. We simulate the natural histories of cancer among $n=40,000$ individuals under the following distribution of first-screen age, inter-screen interval and age of last follow-up. The age of the first screen is sampled from the categorical distribution with weights $w_{k}\propto\exp\{-k/5\}$ for $k=40,41,\dots,80$. Inter-screen intervals are sampled from the distribution $\textsf{Poisson}(0.5)$ plus $1$ year, resulting in a mean inter-screen interval of $1.5$ years. Finally, for individual $i=1,\dots,n$, the age of last follow-up is $\min\{100,t^{i}_{1}+W_{i}\}$ with $W_{i}\sim\textsf{Exp}(1/5)$ independently over $i$, implying that individuals remain in the screening program for $5$ years on average.
We use the independent prior distribution $\pi(\lambda_{\textsf{H}},\lambda_{\textsf{prog}},\psi,\beta)=\pi(\lambda_{\textsf{H}})\pi(\lambda_{\textsf{prog}})\pi(\psi)\pi(\beta)$ with
$\pi(\lambda_{\textsf{H}})=\textsf{Ga}(1,0.01),\quad\pi(\lambda_{\textsf{prog}})=\textsf{Ga}(1,0.01),\quad\pi(\psi)=\textsf{Be}(1,1),\quad\pi(\beta)=\textsf{Be}(38.5,5.8),$
where $\textsf{Ga}(a,b)$ corresponds to the gamma distribution with density proportional to $x^{a-1}\exp\{-bx\}$. We choose these distributions for their flexibility and, in the case of $\pi(\beta)$, for its conjugacy (Theorem 3.1). The prior distribution on $\lambda_{\textsf{H}}$, $\lambda_{\textsf{prog}}$ and $\psi$ are weakly informative, and that of $\beta$ places $95\%$ of its mass on the interval $(0.76,0.95)$, reflecting prior empirical data on the performance of mammography screening 32.
Next, we apply our MCMC sampler to the simulated dataset and estimate the values of the parameters $\lambda_{\textsf{H}}$, $\lambda_{\textsf{prog}}$, $\psi$ and $\beta$. We independently run $20$ MCMC samplers initialized at over-dispersed values of these parameters (the latent variables need not be initialized, see Section 4). The initial values $\lambda_{\textsf{H}}^{(0,m)}$, $\lambda_{\textsf{prog}}^{(0,m)}$, $\psi^{(0,m)}$ and $\beta^{(0,m)}$ of the $m$th sampler $(m=1\dots,20)$ are sampled from a wide interval around the true values, mutually independently,
$10^{4}\times\lambda_{\textsf{H}}^{(0,m)}\sim\textsf{U}(0.2,2),\quad\lambda_{\textsf{prog}}^{(0,m)}\sim\textsf{U}(0.01,0.1),\quad\psi^{(0,m)}\sim\textsf{U}(0,1),\quad\beta^{(0,m)}\sim\textsf{U}(0.7,0.95).$
For each chain, we set the warm-up period to $10^{4}$ iterations and run the sampler for an additional $2\times 10^{4}$ iterations. For each parameter, the potential scale reduction factor, computed on the post warm-up thinned draws with the R package coda 33, is smaller than $1.005$ and the effective sample size is superior to 2,500, indicating that the chains—which were initialized at over-dispersed values—converged during the warmup phase and mixed well once they reached stationarity 30. This is a remarkable achievement given the high dimensional latent space of 80,000 latent variables.
In addition to exploring the parameter space efficiently, the sampler also explores the space of the latent variables swiftly. Figure 4 displays the MCMC draws of the latent variables for an individual whose cancer was screen-detected at age $58$ years, after a series of five consecutive negative screens. The individual had $5$ negative screens at ages $46$, $48$, $51$, $53$ and $55$. The traceplots of the indolent indicator (Figure 4(a)) and the pre-clinical transition time (Figure 4(b)) indicate that the sampler rapidly explores the latent space. In particular, the frequent switches between the values 0 and $1$ of the binary draws in Figure 4(a) show that the Markov chain is not stuck in any particular configuration of the latent variables for any large number of iterations.
An interesting feature of our data augmentation approach is that we obtain a patient-specific estimate of the onset time of pre-clinical cancer, a quantity unobservable in practice. Figure 4(c) illustrates the MCMC approximation of the marginal posterior distribution $\pi\left(\tau^{HP}\mid\mathbf{Y}\right)$ for the above individual with a screen-detected cancer at age 58 years. Because the screening sensitivity is estimated to be between $0.75$ and $0.90$, it is most likely that the time of onset took place after the last screen at $55$ years, less likely that it did between the screens at ages $53$ and $55$, and very unlikely before the screen at age $53$. As illustrated by the discontinuities in this histogram, the screens themselves contribute most of the information available about the unobserved onset age of pre-clinical cancer.
Figure 3: Estimate of the density of the posterior distribution (solid black line) based on the MCMC draws and density of the prior distribution (dashed blue line) for the simulated data.
((a)) Traceplot of $Z^{I}$.
((b)) Traceplot of $Z^{HP}$.
((c)) Histogram of $Z^{HP}$.
Figure 4: MCMC draws from the first chain ($m=1$) for the latent variables of an individual with a screen-detected cancer at age $58$ and five negative screens at ages $(46,48,51,53,55)$. Panel (a) is a traceplot of the latent indolent indicator over an interval of $5.000$ iterations for clarity. Panel (b) is a traceplot of the latent pre-clinical transition age. Panel (c) shows the marginal posterior distribution of the latent pre-clinical transition age; vertical lines depict the ages at which the screens occur.
6 Application to data from the Breast Cancer Surveillance Consortium
We illustrate the utility of our semi-Markov model and accompanying DA-MCMC sampler with an analysis of the screening and cancer diagnosis histories from participants in the Breast Cancer Surveillance Consortium (BCSC), a large population-based mammography screening registry in the US 34. The BCSC registries collect mammography results and link them to breast cancer outcomes and vital status through linkage with regional population-based Surveillance, Epidemiology, and End Results (SEER) programs, state tumor registries, and state death records.
We use the analytic cohort previously described in 11, which includes $35,986$ women aged $50$ to $74$ years who received their first screening mammogram at a BCSC facility between $2000$ and $2018$. Details are provided in Wed Appendix G. Each BCSC registry and the Statistical Coordinating Center (SCC) have received institutional review board approval for all study procedures, including passive permission (one registry), a waiver of consent (six registries), or both depending on facility (one registry), to enroll participants, link data, and perform analytic studies. All procedures are Health Insurance Portability and Accountability Act (HIPAA) compliant. All registries and the SCC have received a Federal Certificate of Confidentiality and other protections for the identities of women, physicians, and facilities who are subjects of this research.
Following 11, we set $t_{0}=30$. We employ the independent prior $\pi(\lambda_{\textsf{H}},\lambda_{\textsf{prog}},\psi,\beta)=\pi(\lambda_{\textsf{H}})\pi(\lambda_{\textsf{prog}})\pi(\psi)\pi(\beta)$ with
where, for a given $\alpha_{\textsf{prog}}$, we choose values for $a_{\textsf{prog}}$ and $b_{\textsf{prog}}$ that induce a weakly informative distribution on the mean sojourn time $\mu_{\textsf{prog}}$ with $0.9$ mass on the interval $(1,9)$ years, reflecting prior knowledge about breast cancer progression. Web Appendix I contains the details.
6.1 Model specification
We consider the $21$ models with $\alpha_{\textsf{H}}\in\{1,1.5,2,2.5,3,4,5\}$ and $\alpha_{\textsf{prog}}\in\{1,1.5,2\}$. We fit each model to the BCSC data and estimate their predictive fit via ALOOCV (see Section 4.2). As no cancer screens in the data are conducted before age $50$, we set $t_{0}=30$. The burn-in period is set to $2\times 10^{4}$ iterations and is followed by an additional $5\times 10^{4}$ iterations; this results in an effective sample size of at least $10^{3}$ for each parameter in every setting.
Figure 5 displays the predictive fit of the models. The relation between the predictive fit and $\alpha_{\textsf{H}}$ follows an inverted-U shape, and, nominally, the best model fit was achieved at $\alpha_{\textsf{H}}=4$ and $\alpha_{\textsf{prog}}=1$. Models with $\alpha_{\textsf{H}}\leq 1.5$ and $2$ of the $3$ models with $\alpha_{\textsf{H}}=2$ have a predictive fit that is significantly worse than the best model at the significance level $0.05$. This suggests that the pre-clinical hazard rate of breast cancer increase at least linearly with age. However, the exact speed of the increase—linear, quadratic, etc—cannot be inferred from these data alone, presumably because the available screens only cover the range 5050-$85$ years, with $97\%$ of them limited to the range 5050-$75$ years.
In contrast, $\alpha_{\textsf{prog}}$ does not have a significant impact on the predictive fit. The three lines in Figure 5, which correspond to the three values of this parameter, are extremely close to one another, and the pairwise differences between the predictive fit of models with the same $\alpha_{\textsf{H}}$ did not reach statistical significance at $\alpha=0.05$. This indicates that the data do not favor any particular value for this parameter. This is to be expected because the endpoints of the pre-clinical sojourn times of progressive cancer are mostly unobserved. Indeed, the start age is never exactly observed; furthermore, as the screens are imperfect, this value is not even interval censored. In addition, the end age of the sojourn is exactly observed only for the interval-detected group ($0.18\%$ of the sample) and is right-censored for the other individuals. This makes the task of learning the exact shape of the distribution of sojourn times from the observed data alone extremely difficult.
We now proceed by choosing among the models whose predictive fit is not significantly worse than the best model the one with values for $\alpha_{\textsf{H}}$ and $\alpha_{\textsf{prog}}$ that we deem most plausible given our medical understanding of the progression of breast cancer. First, we choose a value of $\alpha_{\textsf{prog}}$ that implies a realistic predictive distribution of sojourn times among progressive cancers. Conditional on $\lambda_{\textsf{prog}}$, the sojourn time of a progressive cancer—denoted by $\sigma_{\textsf{prog}}$—follows the Weibull distribution $\textsf{Wei}\left(\lambda_{\textsf{prog}},\alpha_{\textsf{prog}}\right)$. Its predictive distribution
$\pi(\sigma_{\textsf{prog}}\mid\mathbf{Y},\alpha_{\textsf{prog}})=\int\textsf{Wei}\left(\sigma_{\textsf{prog}};\lambda_{\textsf{prog}},\alpha_{\textsf{prog}}\right)\pi\left(\lambda_{\textsf{prog}}\mid\mathbf{Y},\alpha_{\textsf{prog}}\right)d\lambda_{\textsf{prog}},$ (17)
is therefore a rate-mixture of Weibull distributions, in which the Weibull rate is integrated out. We approximate the posterior distribution $\pi\left(\lambda_{\textsf{prog}}\mid\mathbf{Y},\alpha_{\textsf{prog}}\right)$ with the empirical distribution of the MCMC draws $\left\{\lambda_{\textsf{prog}}^{(1)},\dots,\lambda_{\textsf{prog}}^{(M)}\right\}$, giving the Rao-Blackwell estimate of the predictive distribution (17) 27
Web Appendix H compares the estimate (18) for the models with $\alpha_{\textsf{H}}=2$ and $\alpha_{\textsf{prog}}\in\{1,1.5,2\}$. Under $\alpha_{\textsf{prog}}=1$, the predictive distribution has a large amount of mass concentrated at 0 and a fat tail. In contrast, under the semi-Markov models with $\alpha_{\textsf{prog}}\in\{1.5,2\}$ the density of the predictive distribution has a thin tail, and no mass concentrated at 0. As a result, the predictive probability of a sojourn time less than $0.5$ year or more than $15$ years is almost $20\%$ under $\alpha_{\textsf{prog}}=1$, while it is only $1.3\%$ under $\alpha_{\textsf{prog}}=2$. A similar patterns is observed across all values of $\alpha_{\textsf{H}}$. We thus opt for $\alpha_{\textsf{prog}}=2$ because it results in a more realistic model of sojourn times.
Next, we choose the value of $\alpha_{\textsf{H}}$. We opt for $\alpha_{\textsf{H}}=2.5$ because it agrees the most with the trend in age-specific data from SEER. Figure 6 presents the posterior distribution of key parameters under the chosen model. The posterior mean of the indolence probability $\psi$ is $0.057$, with $95\%$ credible interval (0.0020.170.002-0.17), and that of the mean sojourn time $\mu_{\textsf{prog}}$ is $5.6$ years ($95\%$ CI: 4.17.14.1-7.1). Finally, the cumulative probability of pre-clinical onset by age 80 years is estimated to be $12.3\%$ ($95\%$ CI: $11.2\%-13.4\%$), which is consistent with an approximate 1 in 7 lifetime risk of developing breast cancer 35.
((a)) Posterior distribution of $\psi$.
((b)) Posterior distribution of $\mu_{\textsf{prog}}$.
((c)) Estimated pre-clinical hazard function.
Figure 6: Results of the analysis of the data set from the Breast Cancer Surveillance Consortium data set. Panels (a) and (b) display the marginal posterior distributions of the fraction of indolent pre-clinical cancers ($\psi$) and the mean sojourn time among progressive pre-clinical cancers ($\mu_{\textsf{prog}}$), respectively. The posterior means (dashed red line) and 9595% credible intervals (dotted red lines) are included. The dashed blue lines correspond to the prior distributions. Panel (c) shows the estimated pre-clinical hazard rate as a function of age, with the posterior median (solid line) and 9595% credible band (dashed lines) indicated.
6.2 Estimation of the overdiagnosis rate
From the MCMC draws, one can estimate the posterior distribution of any function of the parameters. Here, we focus on the extent of overdiagnosis in a screening program, defined as the the mammographic detection of cancer that would not become clinically evident in the woman’s remaining lifetime. There are two contributions to overdiagnosis. First, by definition, every indolent cancer detected on mammography is overdiagnosed. Second, progressive screen-detected cancer are overdiagnosed if the time from screen-detection to death from a breast cancer-unrelated cause is shorter than the time from screen-detection to clinical progression (also referred to as the lead-time).
We use the breast cancer natural history estimates from the BCSC cohort together with life tables that exclude death from breast cancer to estimate overdiagnosis rates, overall and by source, in a biennial screening program from age $50$ to $74$ years; details of the calculation are provided in web Appendix J. Table 1 provides a side-by-side comparison of the three models with $\alpha_{\textsf{H}}=2.5$. The three models yield similar estimates for the overdiagnosis rate and the respective contributions from indolence and mortality from other sources, showing that these results are robust to the choice of $\alpha_{\textsf{prog}}$. This is reassuring given that the value of this parameter could not be learned from the data. The overdiagnosis estimates are similarly robust to values of $\alpha_{\textsf{H}}$ around $2.5$. Moreover, the estimates we obtain are in line with those in 11, which are based on a Markov model with a piece-wise constant onset rate for pre-clinical cancer, corroborating the claim that approximately 1 in 7 screen-detected breast cancers are overdiagnosed 11.
Table 1: Estimates of the overdiagnosis rate and the respective contribution of indolence and mortality from other causes for the nine models with $(\alpha_{\textsf{H}},\alpha_{\textsf{prog}})=\{2,2.5,3\}\times\{1,1.5,2\}$ and the model in 11. The posterior means and $95\%$ credible intervals are reported.
$\mathbf{\alpha_{\textsf{H}}}$ $\mathbf{\alpha_{\textsf{prog}}}$ Overdiagnosis rate (%) Source Mean sojourn time Probability of indolencea $(\%)$
Indolence (%) Mortality from other causes (%)
22 11 $18.8$ $(6.5,35.9)$ $8.1$ $(0.2,28.0)$ $10.7$ $(2.8,23.0)$ $6.6$ $(5.0,8.7)$ $4.3$ $(0.1,13.1)$
22 $1.5$ $16.9$ $(7.9,36.5)$ $12.7$ $(3.8,35.7)$ $4.1$ $(0.7,10.0)$ $5.6$ $(4.0,7.1)$ $5.7$ $(0.2,15.6)$
22 22 $16.9$ $(5.6,33.1)$ $11.7$ $(1.0,31.4)$ $5.2$ $(0.9,13.2)$ $5.1$ $(3.4,7.1)$ $9.0$ $(0.5,21.8)$
$2.5$ 11 $14.0$ $(4.1,26.5)$ $7.3$ $(0.2,21.7)$ $6.7$ $(1.4,16.2)$ $7.0$ $(5.3,9.0)$ $4.3$ $(0.1,14.8)$
$2.5$ $1.5$ $13.2$ $(3.9,26.0)$ $5.9$ $(0.2,20.0)$ $7.3$ $(1.6,17.1)$ $5.8$ $(4.4,7.4)$ $4.8$ $(0.2,15.5)$
$2.5$ 22 $12.2$ $(3.6,23.8)$ $4.7$ $(0.1,16.4)$ $7.5$ $(1.8,17.5)$ $5.6$ $(4.1,7.3)$ $5.8$ $(0.1,17.8)$
33 11 $11.8$ $(3.4,23.6)$ $4.5$ $(0.1,15.5)$ $7.3$ $(1.8,16.8)$ $7.2$ $(5.5,9.2)$ $4.3$ $(0.1,14.7)$
33 $1.5$ $12.0$ $(3.2,25.4)$ $5.1$ $(0.1,19.6)$ $6.9$ $(1.6,16.0)$ $6.0$ $(4.6,7.5)$ $4.0$ $(0.1,13.0)$
33 22 $18.9$ $(9.4,42.1)$ $15.2$ $(5.4,41.5)$ $3.8$ $(0.6,9.2)$ $5.7$ $(4.3,7.0)$ $4.8$ $(0.2,16.4)$
Model in11 $15.4$ $(9.4,26.5)$ $6.1$ $(0.2,20.1)$ $9.3$ $(5.5,13.5)$ $6.6$ $(4.9,8.6)$ $4.5$ $(0.1,14.8)$
{tablenotes}
Proportion of pre-clinical cancers that are indolent.
7 Discussion
This paper introduces a data-augmented MCMC sampler for fitting semi-Markov mixture models of cancer natural history to individual-level screening and cancer diagnosis histories. By allowing for age-dependent pre-clinical onset hazards, non-exponential sojourn times, and a mixture of indolent and progressive tumors, we substantially expand the range of cancer progression models and dynamics amenable to rigorous Bayesian inference. These advances mark a departure from prior approaches relying on simple Markov models for tractability. Our sampler efficiently explores the high-dimensional joint space of model parameters and latent variables, and automatically quantifies individual-level uncertainty of clinically relevant quantities such as a person’s pre-clinical onset time, indolence status, and sojourn time.
A key application of our methodology is the estimation of screening-related overdiagnosis rates. In the motivating example of breast cancer, overdiagnosis has been documented as the chief harm of screening, yet there remains substantial uncertainty about its true extent. Due to the counterfactual nature of overdiagnosis, its estimation is methodologically challenging and, as in our case of observational data from a screened cohort, relies on sound estimates of the latent tumor progression dynamics. Applying our method to a high-quality cohort of women undergoing mammography screening, we make two important observations. First, our breast cancer overdiagnosis estimates are similar to those previously derived from the same cohort using a slightly different model structure, thus adding weight to the finding that approximately 1 in 7 screen-detected breast cancers are overdiagnosed. Second, although the data contain insufficient information to delineate the best-fitting sojourn time parameterization, we find that the overdiagnosis estimates were robust to changes in this parameterization. Robustness is particularly important because model misspecification is inevitable when modeling complex biological processes such as cancer progression. Finally, the latent variables provide individual-level estimates of pre-clinical cancer onset times. These data could, for example, be leveraged to derive individualized estimation of overdiagnosis.
The methodology presented here has focused on newly enabling exact inference beyond the Markovian setting. While we show that this increased model flexibility remains tractable for inference in large datasets, we note several limitations that may continue to be addressed in future work. Age-dependent extensions will allow for remaining restrictive assumptions to be relaxed. For example, allowing the sojourn time $\sigma_{P}$ to depend on the age of pre-clinical onset is more consistent with clinical knowledge that cancers with later onset tend to progress more slowly. While the method in principle can tackle semi-Markovian models in broad generality, its advantages are in part hampered by fixing a subset of the parameters. Our empirical results via ALOOCV suggest that incorporating parameters such as the onset shape parameter $\alpha_{H}$ into the DA-MCMC sampler may be practically identifiable in some data settings. Similarly, including patient-level covariates into model will further improve model fit and realism, and may find use toward personalized prevention strategies. We note that likelihood-free approaches are readily applicable to include such features and more complex parametrizations, albeit at the cost of intensive simulation-based computation. Structured parametrizations such as log-linear relationships may remain amenable to an exact sampling approach, and have also been studied using EM algorithms 36. Future extensions of the method and software targeting such covariate-dependent rates, as well as nonparametric specifications of the hazard, are promising directions to expand the contributions in this article.